Differentiate between half adder and full adder? Support your answer by presenting the truth table for each.
Answer #1

Answer:

Half adder and Full adder both are belongs to combinational logic circuits.

1.In Half adder addition of 2 one bit digits is performed when come to full adder it adds three one bit digits. ex: half-adder 1 + 0 = 1 (sum) . full-adder 1 + 0 + 0 = 1 (sum).

2.The one of the difference between half adder and full adder is half adder contains one EX_OR and one AND Gate but when come to full adder it contains two EX-OR gates and two AND gates.

3.This two adders has 2 outputs.one is sum and another one is carry.

4.Half-adder doesn't take carry from previous addition when come to full-adder it will happen.

5.we can design a full-adder using the 2 half-adders.

Half-adder :

sum = x ⊕ y and carry = x.y

Full-adder :

sum = x ⊕ y ⊕ c and carry = x.y + c(x⊕y)